OpenStudy (deadshot):

What is the 19th term of a geometric sequence where \[ a_1 = 14\] and \[a_9 = 358.80\]? Round the common ratio and 19th term to the nearest hundredth.

OpenStudy (campbell_st):

well if the series is geometric you know \[a_{9} = a_{1} \times r^{9 -1}\] which can be used to find the common ratio. \[358.8 = 14 \times r^8\] so you need to solve for r, the common ratio \[r = \sqrt[8]{\frac{358.8}{14}}\] when you have r you can go and find the 19th term \[a_{19} = a_{1} \times r^{19 -1}\] hope this helps.